OpenStudy (anonymous):

\[x ^{2}-4x+3+=0\] SOMEONE HELP ... QUADRATIC FORMULA \[x ^{2}-4x+3+=0\] SOMEONE HELP ... QUADRATIC FORMULA @Mathematics

OpenStudy (lilg132):

-3 -1 (x - 3)(x-1) x^2 - x -3x +3 = 0 x - 3 = 0 x - 1 = 0 x = 0 +3 x = 0 + 1 x = 3 x = 1